OpenStudy (anonymous):

Find the equation of the locus of a point which moves so that its distances from the y axis is equal to the square of its distance from (h,0)

OpenStudy (anonymous):

ok how do i begin this question?.. idk

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@freckles

OpenStudy (amistre64):

how do we define distance?

OpenStudy (amistre64):

take a point (x,y) what is its distance from the y axis? what is its distance from (h,0) ?? equate them using the relationship criteria

OpenStudy (anonymous):



OpenStudy (anonymous):

Distance (x,y) to y axis = Distance (x,y) to (0,y) = x Distance (x,y) to (h,0) = √ ( x - h)^2 + (y-0)^2 Directions say : Distance (x,y) to y axis = square of distance from (x,y) to (h,0) x = [√ ( (x-h)^2 + (y-0)^2) ]^2 x = (x-h)^2 + y^2 (x-h)^2 -x + y^2 = 0 This is starting to look like an equation of a circle. Complete the square x^2 -2xh +h^2 -x + y^2 = 0 x^2 -2xh -x + y^2 = -h^2 x^2 -( 2h +1)x + y^2 = -h^2 ( x - (2h+1)/2 ) ^2 + y^2 = -h^2 + ((2h+1)/2 ) ^2 ( x - (2h+1)/2 ) ^2 + y^2 = h + 1/4 What do we know about this equation . The locus of points is a circle with center ( (2h+1)/2, 0 ) and a radius √ (h + 1/4) test this is true, plug in h = 6 for example , and graph it using a graphing utility or by hand I would recommend Desmos calculator, as it lets you use a slider to check different values of h https://www.desmos.com/calculator/ftm9ndy1vj Verifying it graphically is like having someone double check your work for free. Note that for h < -1/4 there is no meaningful output, can you see why?